### v3

# REINFORCE++: An Efficient RLHF Algorithm with Robustness to Both Prompt and Reward Models

Abstract
Reinforcement Learning from Human Feedback (RLHF) is crucial in aligning large language models (LLMs) with human values and preferences. While state-of-the-art applications like ChatGPT/GPT-4 commonly employ Proximal Policy Optimization (PPO), including a critic network introduces significant computational overhead. REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network. However, these approaches face challenges in accurate advantage estimation. Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on more straightforward prompts and vulnerability to reward hacking. To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the normalized reward of a batch as the baseline. Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation. Furthermore, it achieves superior generalization in RLHF and long chain-of-thought (CoT) settings compared to REINFORCE-based methods. The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning large language models (LLMs) with human values and preferences  (Vemprala et al., 2023; Achiam et al., 2023; Ouyang et al., 2022a; Shen & Zhang, 2024). Despite the emergence of non-RL alternatives like DPO  (Rafailov et al., 2023), state-of-the-art applications such as ChatGPT/GPT-4  (Vemprala et al., 2023; OpenAI, 2023), Claude  (Anthropic, 2023), and Gemini  (Team et al., 2023) continue to rely on RL algorithms, particularly PPO, for policy optimization. However, PPO  (Schulman et al., 2017) requires a critic network, introducing substantial computational overhead and memory demands that limit large model alignment in small-scale clusters. To address this, researchers have proposed various REINFORCE-based methods that eliminate the critic network, including ReMax  (Li et al., 2023), REINFORCE Leave One-Out (RLOO)  (Wu et al., 2024), and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024). Furthermore, Deepseek-R1 demonstrates the effectiveness of the REINFORCE-based method in the long-form CoT setting  (Guo et al., 2025), which achieved state-of-the-art performance on challenging datasets using GRPO with rule-based rewards.
Without the critic network, REINFORCE-based methods often struggle to accurately estimate individual tokens’ advantages. Various REINFORCE-baseline approaches have been proposed to address this limitation, but each has significant drawbacks. ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation. RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO utilizes a normalized reward across all responses. These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking. Furthermore, these methods calculate reward baselines separately for each prompt, leading to overfitting and instability on specific training prompts during optimization. Therefore, these methods require carefully curating prompt sets particular to each task.
To address these challenges, we propose REINFORCE++, a novel REINFORCE-based method that eliminates the critic model from PPO and uses the mean reward of a global batch as the baseline. This approach prevents overfitting to specific training prompts and demonstrates robustness across both Bradley-Terry and rule-based reward models. Notably, REINFORCE++ eliminates the need for prompt set truncation and achieves strong generalization performance in both RLHF and long CoT RL settings.
The remainder of this paper is organized as follows. In Section 2, we analyze the differences between various REINFORCE-based RLHF methods, demonstrating the ineffectiveness of prompt-specific reward baselines and the overfitting problems in RLOO and GRPO methods. Section 3 presents REINFORCE++, detailing its implementation and advantages. In Section 4, we evaluate REINFORCE++ through comprehensive experiments. We first compare REINFORCE++ with other RLHF methods using both Bradley-Terry and Rule-Based Reward Models, showing its superior or comparable performance. Then, under the long-form Chain-of-Thought (CoT) setting, we focus on comparing REINFORCE++ with GRPO, the current state-of-the-art REINFORCE-based method, due to computational constraints. Through experiments on a curated small set of carefully selected prompts, we highlight GRPO’s weakness in overfitting specific prompts. Our subsequent experiments on long-form CoT tasks demonstrate REINFORCE++’s superior performance on challenging test datasets, indicating better Out-of-Distribution (OOD) generalization. In Section 5, we conclude our method and future work directions.
Figure 1: The comparison of PPO, Remax, GRPO, RLOO, and REINFORCE++.
2 Background and Related Work
State-of-the-art applications such as ChatGPT/GPT-4, Claude, and Gemini utilize reinforcement learning algorithms, such as PPO, for policy optimization. In particular, PPO optimizes LLMs by maximizing the following surrogate objective:
$\mathcal{L}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(O|q)}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}\min\left(s_{t}(\theta)A_{t},\ \text{clip}(s_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$ (1)
where:
$s_{t}(\theta)=\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}$ (2)
PPO is an actor-critic-based reinforcement learning (RL) algorithm where the critic model consumes significant training resources. Consequently, many researchers have proposed a series of REINFORCE-based methods, such as ReMax, RLOO, and GRPO, to avoid the computational overhead associated with the critic model while still obtaining relatively accurate token-wise advantage estimations. These methods design alternative techniques to calculate the baseline reward for each prompt as the advantage estimation.
Specifically, the advantage is computed to evaluate the performance of an action relative to a baseline in PPO. It starts by collecting samples of states, actions, rewards, and the following states. The advantage is then calculated using Generalized Advantage Estimation (GAE), which combines the temporal difference error $\delta_{q,o_{t}}=r_{t}+\gamma V(o_{t+1})-V(o_{t})$ over a series of time steps:
$A_{q,o_{t}}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}$ (3)
where $\lambda$ is a parameter balancing bias and variance. To obtain an accurate advantage estimation and generalize to unseen tokens, PPO must use the critic model to learn the advantage function, obtaining an accurate advantage estimation and generalizing to unseen tokens.
As shown in Figure 1, ReMax adopts the greedy decoding method to generate a response and obtain its reward as the baseline reward for this prompt. Accordingly, the advantage of a query $q$ is given by
$A_{q,o_{t}}=r(o_{1:t},q)-r(\hat{o}_{1:t},q)$ (4)
where:
$\hat{o}_{1:t}=\operatorname*{argmax}_{o^{\prime}_{1},\dots,o^{\prime}_{t}}\prod_{i=1}^{t}\pi_{\theta}(o^{\prime}_{i}\mid q,o^{\prime}_{<i})$ (5)
Notably, ReMax does not use this response to train the model.
Both RLOO and GRPO sample multiple responses for a prompt. RLOO adopts the average rewards of all other samples for the current prompt as a baseline, which is shown in Equation 6.
$A_{q,o_{t}^{(i)}}=r(o_{1:t}^{(i)},q)-\frac{1}{k-1}\sum_{j\neq i}r(o_{1:t}^{(j)},q)$ (6)
GRPO, on the other hand, adopts the group relative advantage estimation method, which uses the mean reward divided by the standard deviation of all sampled responses for the current prompt as the baseline reward, which is shown in Equation 7.
$A_{q,o_{t}^{(i)}}=\frac{r(o_{1:t}^{(i)},q)-\text{mean}\bigl{(}\{r(o_{1:t}^{(j)},q)\}_{j=1}^{k}\bigr{)}}{\text{std}\bigl{(}\{r(o_{1:t}^{(j)},q)\}_{j=1}^{k}\bigr{)}}$ (7)
However, given a diverse dataset, we consider that the baseline reward for each prompt is not essential in the RLHF setting. Although the baseline reward of each prompt can provide relatively accurate advantage estimation for each training prompt and help the model learn the response with the highest reward under a prompt, it also accelerates the reward hacking and overfitting problems. It is worth noting that RLHF has two key differences from traditional RL problems:
  • In traditional RL problems, we train an RL policy and test it in the same environment. However, the RLHF method trains a prompt set and tests on another dataset, even an out-of-distribution (OOD) dataset.
  • In traditional RL problems, there is always a golden reward. In contrast, the RLHF method always uses a reward model or rule-based reward, which may encounter reward hacking.
Accordingly, both reward hacking and overfitting problems may degrade model performance. Specifically, optimizing multiple responses for a prompt in a batch using methods like RLOO and GRPO tends to overfit the best response under specific simple prompts, ultimately degrading the model’s generalization. Furthermore, optimizing multiple responses for a prompt within a single batch reduces the diversity of the model’s outputs, leading to a low diversity in token-level advantage distribution and causing overfitting on those tokens. In contrast, PPO does not suffer significantly from this issue, as the value network continues to be trained and retains the learned advantages, thereby supporting a more generalizable token-wise advantage.
Accordingly, to avoid overfitting specific prompts and increase prompt diversity in a training batch, REINFORCE++ can sample one response for each prompt and normalize the token-wise advantage in the global batch size to increase training stability.
3 Method
REINFORCE++ optimizes the PPO objective and employs the clipping strategy defined in Equation 1. To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward. Accordingly, the advantage of REINFORCE++ is the normalized reward-to-go in reinforcement learning:
$A_{q,o_{t}}=r(o_{1:t},q)-\beta\cdot\sum_{i=t}^{T}\text{KL}(i)$ (8)
where:
$\text{KL}(t)=\log\left(\frac{\pi_{\theta_{\text{old}}}^{\text{RL}}(o_{t}|q,o_{<t})}{\pi^{\text{SFT}}(o_{t}|q,o_{<t})}\right)$ (9)
The gradient of the token-level KL penalty has been theoretically proven to be unbiased concerning the $k_{3}$ loss of GRPO in RLHF. Additionally, we normalize this advantage across the global batch for all prompts:
$A_{q,o_{t}}^{\text{norm}}=\frac{A_{q,o_{t}}-\text{mean}\left(A_{q,o_{t}}\right)}{\text{std}\left(A_{q,o_{t}}\right)},$ (10)
REINFORCE++ removes the prompt-specific reward baseline and does not optimize multiple responses from a single prompt within a training batch. Compared to PPO, REINFORCE++ eliminates the critic model and sets the discount factor to $1$. The detailed implementation of our algorithm is provided in Algorithm 1, with further algorithmic details discussed therein.
Algorithm 1 REINFORCE++
1:Initial policy model $\pi_{\text{init}}$, reward models $R$, task prompts $\mathcal{D}$
2:policy model $\pi_{\theta}\leftarrow\pi_{\text{init}}$
3:for step $=1,\ldots,M$ do
4:     Sample a batch $\mathcal{D}_{b}$ from $\mathcal{D}$
5:     Update the old policy model $\pi_{\text{old}}\leftarrow\pi_{\theta}$
6:     Sample one output $o\sim\pi_{\text{old}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{b}$
7:     Compute rewards $r_{i}$ for each sampled in a batch of output $o_{i}$ by running $R$
8:     Compute $A_{q,o_{t}}^{norm}$ for the $t$-th token of $o$ for prompt $q$ through Equation 8 to 10
9:     for iteration $=1,\ldots,t$ do
10:         Update the policy model $\pi_{\theta}$ by maximizing the REINFORCE++ objective (Equation 1)
11:     end for
12:end for
13:$\pi_{\theta}$
4 Experiments
The empirical evaluation of REINFORCE++ was conducted across diverse test scenarios to assess its performance comprehensively. In experiments using the Bradley-Terry Reward Model, we thoroughly compare REINFORCE++ against existing methods, including ReMax, RLOO, GRPO, and PPO. Due to computational constraints, we focused our comparisons between REINFORCE++ and GRPO, the current state-of-the-art REINFORCE-based method for experiments in the long-form CoT setting.
4.1 Performance with Outcome Reward Model
Experimental Setup
This study investigates reinforcement learning with human feedback (RLHF) in a general-domain setting. The experimental process began with an instruction-following policy model, fine-tuned on diverse datasets reflecting general-domain language tasks. The model was subsequently refined using reinforcement learning algorithms guided by a reward model. Specifically, a Bradley-Terry reward model was trained on human-generated preference data obtained through pairwise comparisons of model outputs. The reward model evaluates the policy model’s responses based on helpfulness, accuracy, coherence, and alignment with human intent and outputs an outcome reward value for each Bai et al. (2022).
Reward Model
Following the approach proposed by Ouyang et al. (2022b), we initialize the reward model using a SFT model  [^1] https://huggingface.co/OpenRLHF/Llama-3-8b-sft-mixture. To adapt the model for preference learning, we replace the final layer with a linear head that produces a scalar output. The reward model is then trained using the negative log-likelihood loss function defined as:
$\mathcal{L}_{\text{RM}}(\theta)=-\mathbb{E}_{\left(q,o^{+},o^{-}\right)\sim\mathcal{D}}\left[\log\sigma\left(r_{\theta}\left(q,o^{+}\right)-r_{\theta}\left(q,o^{-}\right)\right)\right]$ (11)
Dataset
For the training of our reward model, we leveraged an extensive dataset comprising approximately 700,000 pairs of human preference data [^2] https://huggingface.co/datasets/hendrydong/preference_700K aggregated from multiple publicly available datasets. These datasets provide a rich diversity of contexts and preferences, facilitating the reward model in effectively capturing nuanced human judgment. To systematically prompt the policy model for response generation, we curated a carefully balanced set of 20,000 prompts sampled from diverse sources [^3] https://huggingface.co/datasets/RLHFlow/prompt-collection-v0.1, ensuring comprehensive coverage of various scenarios and domains. This diversity promotes robustness and generalization in the responses generated by the policy model.
Score Length Per Token Score
REINFORCE++ 46.7 832 0.0561
REINFORCE++ Baseline 44.2 834 0.0530
GRPO 46.8 860 0.0544
RLOO 44.6 866 0.0515
ReMax 45.1 805 0.0560
Table 1: Comparison between GRPO and REINFORCE++ on Score and Length. Better results for each reward model are highlighted in bold.
Experimental Results
We use Chat-Arena-Hard Li et al. (2024) to evaluate our models. As illustrated in Table 1, GRPO achieves a slightly superior overall score of 46.8 compared to 46.7 for REINFORCE++. However, GRPO generates longer sequences of an average of 860 tokens, whereas REINFORCE++ produces shorter outputs of only 832 tokens. Consequently, when evaluating performance on a per-token basis, REINFORCE++ outperforms GRPO, achieving a higher per-token score of 0.0561 compared to GRPO’s 0.0544. The table suggests that REINFORCE++ provides more efficient outputs despite the preference bias of the evaluation model on the length.
Figure 2: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and KL Divergence with Outcome Reward Model.
Results Analysis
We plot the test reward and KL divergence curves in Figure 2 to provide a clear insight into the comparative behavior of GRPO and REINFORCE++. Initially, GRPO achieves significantly higher rewards and outperforms REINFORCE++ throughout training. However, closer inspection reveals that the superior rewards achieved by GRPO are mainly due to reward hacking. Specifically, the rapid increase in KL divergence of GRPO suggests that the model is hacking the reward signal rather than improving its generalization performance, resulting in inflated reward values without corresponding gains on the test set. In contrast, REINFORCE++ shows a more gradual but stable increase in reward, accompanied by a modest rise in KL divergence. The result indicates a favorable trade-off, with substantial improvements in the reward being achieved with minimal deviation from the reference model. The comparatively small increase in KL divergence for REINFORCE++ highlights that each divergence unit is translated into a more effective and robust policy improvement, suggesting a higher KL-to-reward conversion efficiency. To substantiate our argument, we evaluated our model on Out-of-Distribution tasks with the setting of RLHF, which primarily includes mathematical problem tasks (GSM8K, MATH) and code generation tasks (HumanEval, MBPP). The evaluation results are presented in Table 2.
GSM8K MATH HumanRval MBPP Avg.
REINFORCE++ 96.21 75.20 85.98 84.39 85.45
REINFORCE++ Baseline 95.98 72.40 78.66 85.45 83.12
GRPO 96.21 73.80 80.49 83.33 83.46
RLOO 96.44 72.40 79.27 82.54 82.67
ReMax 96.59 75.40 78.66 82.54 84.05
Table 2: Comparison between different advantage estimate methods on OOD benchmarks. Better results for each reward model are highlighted in bold.
4.2 Performance on Long-form Chain-of-Thought Tasks
4.2.1 Analysis on Small-Scale Datasets
Experiment Setup
We trained the Qwen2.5-Math-7B pre-trained model using only 30 questions and answers from AIME-24 and evaluated its performance on the AIME-25 dataset. Under this limited training setting, GRPO notably demonstrated its weakness by overfitting the small training dataset.
AIME-24 AIME-25
Pass@N N = 1 N = 1 N = 16
GRPO 95.0 0.0 0.4
REINFORCE++ 71.0 2.5 40.0
Table 3: Comparison between GRPO and REINFORCE++ on both training and test datasets (AIME-24 and AIME-25). Better results for each reward model are highlighted in bold.
Experimental Results
As shown in Table 3, GRPO achieves nearly perfect scores (approximately 100) on the training dataset (AIME-24). However, it performs poorly on the testing dataset (AIME-25), scoring almost 0 in both Pass@1 and Pass@16 test settings. In contrast, while REINFORCE++ achieves a more modest score of 71.0 on AIME-24, it demonstrates better generalization with scores of 2.5 and 40.0 on Pass@1 and Pass@16 test settings, respectively.
Figure 3: Training curves for 15 randomly selected questions during reinforcement learning from zero on a small prompt dataset. Left: Scores trained using GRPO; Right: Scores trained using REINFORCE++
Results Analysis
Furthermore, we analyze the training curves of 15 randomly selected cases. Our findings reveal that GRPO rapidly achieves 100% accuracy (represented by a 1.0 score in Figure 3) within a few steps. In contrast, REINFORCE++ demonstrates a more gradual improvement, typically reaching the same level of accuracy throughout 10 to 20 steps. These observations suggest that GRPO overfits the training set. Further analysis reveals that responses from the GRPO-trained model are significantly shorter, averaging only 30 tokens, compared to 425 tokens for the REINFORCE++-trained model. This evidence confirms that GRPO is more susceptible to overfitting training prompts than REINFORCE++, resulting in poorer generalization performance on test datasets.
4.2.2 RL from Supervised Fine-tuned Model
Experimental Setup
In real-world tasks, user prompts exhibit enormous form and intent, making it difficult to systematically assess a given model’s strengths and weaknesses. Using synthetic datasets, we can systematically manipulate key factors such as length and difficulty, allowing for a more direct and interpretable evaluation and analysis of model performance.
Dataset & Hyper-Parameter
Following Logic-RL (Xie et al., 2025), we incorporate the Knights and Knaves (K&K) puzzles (Xie et al., 2024) into RL training as an algorithmically generated dataset for logical reasoning. In these puzzles, each character is either a knight, who always tells the truth, or a knave, who always lies. The objective is to determine each character’s identity based on their statements. A key feature of this dataset is its strong controllability. The length of the prompt is proportional to the number of roles, and the difficulty can be adjusted by modifying the complexity of logical operations. These puzzles serve as unseen data for the original model, making them ideal for evaluating generalization capabilities. Since the model’s performance in logical reasoning depends on its ability to follow instructions and understand context, we did not begin our experiment with the base model. Instead, we selected a model, Qwen2.5-7B-Instruct-1M, with enhanced capability for handling more extended contexts and following control instructions. We keep all hyper-parameters the same and compare reinforcement learning algorithms by varying the advantage estimation function.
Figure 4: Comparison between GRPO and REINFORCE++ on logic benchmarks with different difficulty levels. The lines show the average performance of the two methods.
Experimental Results
Figure 4 presents a comparative analysis of GRPO and REINFORCE++ on the test datasets. Increasing the number of people - a proxy for task difficulty - leads to decreased performance for both methods. In particular, GRPO achieves slightly higher scores in simpler scenarios involving two or three people, but its performance deteriorates significantly as the number of people increases. On the other hand, REINFORCE++ shows more excellent performance stability and achieves better results in scenarios involving four or more people. This trend is particularly evident in the Out-of-Distribution scenario of eight people, a setting not present in the training data, where GRPO scores 20 and REINFORCE++ scores 36. Overall, REINFORCE++ achieves an average score of 62.1, outperforming the average of 55.7. These results indicate that REINFORCE++ generalizes more effectively in complex and OOD scenarios, underscoring its superior robustness compared to GRPO.
Results Analysis
We analyze the training curves in Figure 5. Our findings indicate that GRPO achieves significant reward values within a few hundred steps. In contrast, REINFORCE++ shows a more moderate and gradual increase in reward, ultimately converging to a higher stable value. Regarding response length, the GRPO-trained model generates considerably shorter responses, averaging approximately 600 tokens, whereas REINFORCE++ consistently produces longer responses of around 1000 tokens. This disparity suggests that GRPO generates shorter, potentially superficial responses, hinting at possible overfitting or memorization rather than meaningful reasoning. Consequently, the model trained with GRPO might generalize poorly on unseen test cases, reinforcing our earlier observations of its susceptibility to overfitting.
4.2.3 RL from Zero Setting
Experimental Setup
Inspired by DeepSeek-R1 (Guo et al., 2025), we apply reinforcement learning with verified reward (RLVR) to a base model for mathematical tasks to evaluate different strategies on the model’s reasoning ability. Previous studies suggest starting from a powerful base model with strong reasoning potential; therefore, we choose Qwen2.5-Math-Base [^4] https://huggingface.co/Qwen/Qwen2.5-Math-7B as the base model.
Figure 5: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and Response Length from Supervised Fine-tuned Model. The $x$ axis represents the training steps.
Dataset
The dataset primarily comprises difficulty levels 3 to 5 of the MATH training split. Due to the context length limitations of the base model, we select approximately 8,000 shorter prompts from the dataset to maximize the observation of variations in the model’s output length.
Hyper-Parameter
We keep all hyper-parameters the same and compare reinforcement learning algorithms by varying the advantage estimation function. In each exploration step, we select 32 questions and generate eight answers per question to maintain a balance between computational efficiency and stability. The actor learning rate is $5\times 10^{-7}$ to optimize the convergence of policy updates. A discount factor ($\gamma$) 1.0 is used to emphasize long-term reward accumulation. The clipping parameter ($\epsilon$) is set to 0.2 to stabilize training and regulate policy updates. The KL penalty coefficient ($\beta$) is set to 0.001 to ensure controlled divergence from the reference model.
AIME-24 AMC-23 MATH-500
Pass@N N = 8 N = 8 N = 1
GRPO 18.96 59.22 73.00
REINFORCE++ 21.04 60.47 72.00
Table 4: Comparison between GRPO and REINFORCE++ on both training and test datasets (AIME-24 and AIME-25). Better results for each reward model are highlighted in bold.
Experimental Results & Analysis
As shown in table 4, REINFORCE++ consistently outperforms GRPO in the out-of-distribution setting and performs similarly in the other setting. Specifically, on the in-distribution test dataset (MATH-500), GRPO achieves a slightly higher Pass@1 score of 73.00 compared to 72.00 for REINFORCE++. However, REINFORCE++ demonstrates superior generalization on the OOD scores, outperforming GRPO on AIME-24 with a Pass@8 score of 21.04 compared to 18.96 and on AIME-25 with a score of 60.47 compared to 59.22. These results indicate that GRPO performs competitively within the training distribution but shows signs of overfitting, whereas REINFORCE++ generalizes better to novel, particularly challenging OOD scenarios.
Figure 6: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and Response Length with Rule Reward Model. The $x$ axis represents the training steps.
5 Conclusion
This paper presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences. Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global batch mean reward as a baseline to prevent overfitting and instability. Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities. Future work includes exploring adaptive normalization techniques, advanced variance reduction methods, and extending REINFORCE++ beyond RLHF settings.
Appendix A Algorithm Details
REINFORCE++-baseline.
For ease of use and reproducibility, we implement our proposed method within the OpenRLHF framework (https://github.com/OpenRLHF/OpenRLHF).
Our experiments above showed that training models using multiple generated responses per prompt can further improve performance. Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation. Specifically, we sample multiple responses for each prompt and compute their average reward as the baseline to reshape the rewards. The advantage for each prompt-response pair is then defined as:
$Adv_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$
$Adv_{q,o_{t}}^{norm}=\frac{Adv_{q,o_{t}}-\text{mean}_{batch}(Adv_{q,o_{t}})}{\text{std}_{batch}(Adv_{q,o_{t}})}$
where $group$ denotes the generated responses corresponding to the same prompt. This baseline calculation is similar to the approach taken by GRPO; however, we deliberately move the standard deviation ($std$) of the group normalization to the global batch normalization. In practice, we observed that including the standard deviation introduces training instability and exacerbates overfitting to easier prompts in the RLHF process. For the KL constraint, REINFORCE++-baseline uses $k_{2}$ instead of the $k_{3}$ divergence approximation method in GRPO. For more details, please refer to the next paragraph.
KL Penalty Design.
In the design of REINFORCE-based RLHF methods, the choice of KL divergence estimation significantly influences algorithmic stability and accuracy. The widely utilized $k_{3}$ estimation in GRPO, defined by the loss function:
$\mathcal{L}_{k_{3}}=\mathbb{E}_{x\sim D,y\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left(\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-\log\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-1\right)$ (12)
Which involve inherent biases due to its approximation nature. Specifically, the gradient of $k_{3}$ estimation approximates the gradient of the $k_{2}$ estimation through linearization:
$k_{2}$ Gradient $:\log x\cdot\nabla_{\theta}\log\pi_{\theta},$ (13)
$k_{3}$ Gradient $:(x-1)\cdot\nabla_{\theta}\log\pi_{\theta},$ (14)
where $x=\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$. Although the $k_{3}$ gradient aligns closely with the $k_{2}$ gradient when the policy is close to the reference ($x\approx 1$), substantial deviations result in two key deficiencies:
  1. 1.
    Bias: As the current policy diverges significantly from the reference, particularly when $\pi_{\text{ref}}\gg\pi_{\theta_{\text{old}}}$, the linear approximation error expands nonlinearly, causing considerable biases.
  2. 2.
    Asymmetry: The function $(x-1)$ exhibits asymmetric behavior depending on whether $\pi_{\theta_{\text{old}}}$ surpasses or falls below $\pi_{\text{ref}}$, further complicating convergence properties.
Therefore, despite its computational simplicity, $k_{3}$ estimation introduces biases and increased variance, suggesting it is not strictly superior to the theoretically unbiased $k_{2}$estimation. Experimental evidence consistently reveals that employing $k_{3}$ estimation in GRPO results in higher variance fluctuations than $k_{2}$ estimation. Further discussion can be found in Liu (2025).
Relationship with PPO.
REINFORCE++ shares certain similarities with PPO in its formulation. Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and employs global batch normalization as the baseline. Mathematically, this relationship is expressed as follows:
$\text{GAE}(\lambda=1,\gamma=1)=\sum_{l=0}^{\infty}r_{t+l}-V(s_{t}),$ (15)
where $V(s_{t})$ denotes the estimated value function at state $s_{t}$. By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term. At the same time, the introduction of global batch normalization further stabilizes training by reducing variance and preventing overfitting to individual prompts.
Other Implementation Details
Mini-Batch Updates
To enhance training efficiency, we implement mini-batch updates with the following characteristics:
  • Batch Processing: Data is processed in smaller, manageable chunks rather than full-batch updates.
  • Multiple Updates: Each mini-batch allows for multiple parameter updates, improving convergence rates.
  • Stochastic Optimization: Introduces beneficial randomness for better generalization.
Reward Normalization and Clipping
We implement comprehensive reward processing to stabilize training:
  • Normalization: Standardizes rewards using z-score normalization to mitigate outliers.
  • Clipping: Constrains reward values within predefined bounds to avoid instability.
  • Scaling: Applies appropriate scaling factors for numerical stability during updates.
Appendix B Acknowledgements
  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the early draft of the paper.
  • Jason Klein Liu: Implemented the experimental code, fine-tuned hyperparameters, wrote the paper, and provided GPU resources.
  • Wei Shen: Led the paper writing and designed and supervised the main experiments.